OpenStudy (anonymous):

Please help! Will fan and medal :) What is the solution to the equation 8 - 2b = -2/3(12b + 6)? b =

OpenStudy (hockeychick23):

Ok so first you need to distribute -2/3 to everything on the right side. So first do -2/3*12b

OpenStudy (anonymous):

@hockeychick23 -8?

OpenStudy (hockeychick23):

yes -8b, now we do -2/3*6

OpenStudy (anonymous):

@hockeychick23 sorry if it takes long my computer is lagging :p -4?

OpenStudy (hockeychick23):

yep -4. so now we have -8b-4=8-2b so now we add 8b to both sides, so we have: 8b-8b-4=8+8b-2b. so now add like terms

OpenStudy (anonymous):

I'm sorry can you "dumb it down" a bit, I'm really bad at math...

OpenStudy (hockeychick23):

ok, so first what's 8b-8b

OpenStudy (anonymous):

0

OpenStudy (hockeychick23):

correct, so we got rid of b on the left side. So now what is 8b-2b?

OpenStudy (anonymous):

6b?

OpenStudy (hockeychick23):

correct, so now we have -4= 6b+8. So now we subtract 8 from both sides so what is -8-4?

OpenStudy (anonymous):

@hockeychick23 -12? Again, sorry if it takes long :p

OpenStudy (hockeychick23):

its cool and correct so we have -12= 6b. So now we divide on both sides so we have: -12/6= ?

OpenStudy (anonymous):

-2?

OpenStudy (hockeychick23):

correct, b=-2

OpenStudy (anonymous):

Yay! Thank you so much, you helped A LOT!

OpenStudy (hockeychick23):

No problem!(:

OpenStudy (anonymous):

@hockeychick23 Sorry to bother but do you think you can check my answer for 1 problem I did?

OpenStudy (hockeychick23):

Sure

OpenStudy (anonymous):

Thank you! Okay brb i'm going to write it down

OpenStudy (anonymous):

 2k-10/-7 =8 k=3?